# SEGAN: semi-supervised learning approach for missing data imputation

### v2

@@ Methods
The source code for all data imputation models originates from the corresponding publications. For all machine learning-based imputation methods, the learning rate is set to 0.3, with the number of iterations set at 100. The MissForest (MissFI) algorithm is configured with 6 decision trees. The Baran algorithm employs AdaBoost as the predictive model. The Multiple Imputation by Chained Equations (MICE) algorithm performs imputation 20 times. For all deep learning-based imputation algorithms, the learning rate is set to 0.001, with a dropout rate of 0.5, training epochs set to 30, batch size at 128, and the ADAM optimizer is used for network training. The Multiple Imputation using Denoising Autoencoders (MIDAE) algorithm consists of 2 fully connected layers with 128 neural units each. The encoder and decoder of the Variational Autoencoder Imputer (VAEI) are fully connected networks, each module having two hidden layers with 20 neurons each, and the latent vector dimensionality is 10. In the Graph Imputation Neural Network (GINN) algorithm, the discriminator is a simple 3-layer feedforward network. In the Generative Adversarial Imputation Nets (GAIN) algorithm, both the generator and discriminator are modeled as 2-layer fully connected networks. The user-specified parameter $\alpha$ for the EDIT-GAIN algorithm is set to 100%. Additionally, in the SEGAN algorithm, the discriminator’s time hint matrix encodes 80% of the mask matrix information, with the default data containing a label rate of 100%.
Result
Table  1 presents the results of missing feature imputation experiments on three incomplete datasets using various data imputation methods. It should be noted that some imputation methods were unable to complete model training within 24 hours, resulting in missing outcomes, which are denoted by ”—”. It is evident that the SEGAN algorithm significantly outperforms all comparison algorithms in terms of imputation performance. In terms of imputation accuracy (i.e., RMSE), the SEGAN algorithm achieves an average improvement of 2.09% over the best comparison algorithm, EDIT, and even a 2.64% increase in imputation accuracy over the CAEGAN algorithm on the Emer dataset. This improvement can be attributed to the SEGAN algorithm’s consideration and analysis of data label information, which enhances the model’s imputation accuracy and stability. It is important to note that due to the high complexity of models such as MissForest (MissFI), XGBoost Imputer (XGBI), Bayesian PCA (Baran), Multiple Imputation by Chained Equations (MICE), Iterative Collaborative Filtering for Recommender Systems (ICLR), DataWig, Robust Recursive Imputation (RRSI), Multiple Imputation using Denoising Autoencoders (MIDAE), and Multiple Imputation with Wasserstein Autoencoders (MIWAE), they could not complete model training within 24 hours. Consequently, these algorithms will not be compared in subsequent experiments.
@@
Figure  2 illustrates the RMSE results of the imputation algorithms as the data missingness rate Rm (i.e., the proportion of missing data in each dataset) increases from 20% to 80%. It is observed that in all cases, the SEGAN algorithm achieves better imputation accuracy, indicated by lower prediction errors (RMSE), compared to the existing optimal imputation algorithms. As the missingness rate Rm increases, the imputation per-formance of the SEGAN algorithm remains more robust.
@@
Fig 2: Comparison of Algorithm Imputation Performance and SEGAN Model Architecture
Furthermore, the imputation accuracy of all algorithms tends to decline with the increase in missingness rate. The reason for this is that as the missingness rate rises, the amount of known data available to the imputation algorithms decreases, leading to a corresponding decline in model imputation accuracy.
Ablation Study Evaluation
In this experiment, we explored the impact of different components within the SEGAN algorithm on its imputation performance. Table 2 presents the RMSE (Root Mean Square Error) for the SEGAN algorithm across various datasets. The S-no-C algorithm refers to SEGAN without the classifier, S-no-D denotes SEGAN without the discriminator, and S-no-R represents SEGAN without the time hint matrix.
@@
Table 2: Ablation Study Evaluation.
@@
Table 3: Comparison of methods on different metrics.
@@
In response to the issue of incomplete data and label information, this paper proposes a semi-supervised learning-based method for missing data imputation named SEGAN, which includes three modules: a generator, a discriminator, and a classifier. The SEGAN algorithm introduces a timing hint matrix to the discriminator to better distinguish between generated and real data and proposes a semi-supervised classifier to address the problem of insufficient data labels. Additionally, this chapter also demonstrates that with the aid of the timing hint matrix and the semi-supervised classifier, the SEGAN algorithm can achieve a unique Nash equilibrium, allowing the generated data distribution and predicted label distribution to converge respectively to the true data distribution and the true label distribution. Ultimately, through extensive experimentation, it is proven that the SEGAN algorithm significantly outperforms existing state-of-the-art methods for multivariate time-series data imputation on both data completion tasks and downstream prediction tasks.
@@ Acknowledgments
This research was funded by the National Key Research and Development Program, China(No. 2023YFC3306303 ),the Major scientific and technological projects of China National Tobacco Corporation(No. 110202201001(JY -01)), and the Zhejiang Provincial “Jianbing”“Lingyan” Re-search and Development Program, China(No. 2023C01213).

### v3

@@ Methods
Implementation details. The source code for all data imputation models originates from the corresponding publications. For all machine learning-based imputation methods, the learning rate is set to 0.3, with the number of iterations set at 100. The MissForest (MissFI) algorithm is configured with 6 decision trees. Baran employs AdaBoost as the predictive model. The Multiple Imputation by Chained Equations (MICE) algorithm performs imputation 20 times. For all deep learning-based imputation algorithms, the learning rate is set to 0.001, with a dropout rate of 0.5, training epochs are set to 30, batch size at 128, and the ADAM optimizer is used for network training. The Multiple Imputation using Denoising Autoencoders (MIDAE) algorithm consists of 2 fully connected layers with 128 neural units each. The encoder and decoder of the Variational Autoencoder Imputer (VAEI) are fully connected networks, each module having two hidden layers with 20 neurons each, and the latent vector dimensionality is 10. In the Graph Imputation Neural Network (GINN) algorithm, the discriminator is a simple 3-layer feedforward network. In the Generative Adversarial Imputation Nets (GAIN) algorithm, both the generator and discriminator are modeled as 2-layer fully connected networks. The user-specified parameter $\alpha$ for the EDIT-GAIN algorithm is set to 100%. Additionally, in the SEGAN algorithm, the discriminator’s time hint matrix encodes 80% of the mask matrix information, with the default data containing a label rate of 100%.
@@
Result
Table  1 presents the results of missing feature imputation experiments on three incomplete datasets using various data imputation methods. It should be noted that some imputation methods were unable to complete model training within 24 hours, resulting in missing outcomes, which are denoted by “--”.
It is evident that the SEGAN algorithm significantly outperforms all comparison algorithms in terms of imputation performance. In terms of imputation accuracy (i.e., RMSE), the SEGAN algorithm achieves an average improvement of 2.09% over the best comparison algorithm, EDIT, and even a 2.64% increase in imputation accuracy over the CAEGAN algorithm on the Emer dataset. This improvement can be attributed to the SEGAN algorithm’s consideration and analysis of data label information, which enhances the model’s imputation accuracy and stability. It is important to note that due to the high complexity of models such as MissForest (MissFI), XGBoost Imputer (XGBI), Bayesian PCA (Baran), Multiple Imputation by Chained Equations (MICE), Iterative Collaborative Filtering for Recommender Systems (ICLR), DataWig, Robust Recursive Imputation (RRSI), Multiple Imputation using Denoising Autoencoders (MIDAE), and Multiple Imputation with Wasserstein Autoencoders (MIWAE), they could not complete model training within 24 hours. Consequently, these algorithms will not be compared in subsequent experiments.
@@
Figure  2 illustrates the RMSE results of the imputation algorithms as the data missingness rate Rm (i.e., the proportion of missing data in each dataset) increases from 20% to 80%. It is observed that in all cases, the SEGAN algorithm achieves better imputation accuracy, indicated by lower prediction errors (RMSE), compared to the existing optimal imputation algorithms. As the missingness rate Rm increases, the imputation performance of the SEGAN algorithm remains more robust.
@@
Fig 2: The prediction performance vs. missing rate $R_{m}$.
Furthermore, the imputation accuracy of all algorithms tends to decline with the increase in the missingness rate. The reason for this is that as the missingness rate rises, the amount of known data available to the imputation algorithms decreases, leading to a corresponding decline in model imputation accuracy.
Ablation Study
In this experiment, we explored the impact of different components within the SEGAN algorithm on its imputation performance. Table 2 presents the RMSE for the SEGAN algorithm across various datasets. The S-no-C algorithm refers to SEGAN without the classifier, S-no-D denotes SEGAN without the discriminator, and S-no-R represents SEGAN without the time hint matrix.
@@
Table 2: The ablation study of SEGAN
@@
Table 3: Post-imputation evaluation on imputation methods
@@
In response to the issue of incomplete data and label information, this paper proposes a semi-supervised learning-based method for missing data imputation named SEGAN, which includes three modules: a generator, a discriminator, and a classifier. The SEGAN algorithm introduces a timing hint matrix to the discriminator to better distinguish between generated and real data and proposes a semi-supervised classifier to address the problem of insufficient data labels. Additionally, this chapter also demonstrates that with the aid of the timing hint matrix and the semi-supervised classifier, the SEGAN algorithm can achieve a unique Nash equilibrium, allowing the generated data distribution and predicted label distribution to converge respectively to the true data distribution and the true label distribution. Ultimately, through extensive experimentation, it is proven that the SEGAN algorithm significantly outperforms existing state-of-the-art methods for multivariate time-series data imputation on both data imputation tasks and downstream prediction tasks.
@@ Acknowledgments
This research was funded by the National Key Research and Development Program, China (No. 2023YFC3306303), the Major scientific and technological projects of China National Tobacco Corporation(No. 110202201001(JY -01)), and the Zhejiang Provincial “Jianbing” Lingyan” Re-search and Development Program, China(No. 2023C01213).
Supported by State Key Laboratory of Massive Personalized Customization System and Technology (H&C-MPC-2023-05-02).